# More than Meets the Eye: Understanding the Effect of Individual Objects on Perceived Visual Privacy

### v1

@@ 4.3. Effect of Environment (H2)
To evaluate H2, we first compared the average comfortability levels across environments. Comfortability level was consistently lower in office settings than in café, with the sharpest drop observed for faces as a foreground PSO (Face: 0.67-0.67; Misc. Paper: 0.37-0.37; Screen: 0.31-0.31). To systematically assess these differences, we applied the Mann–Whitney U test. As shown in Table 5, participants reported significantly lower comfortability (i.e., higher perceived privacy level, PPL) in office contexts across multiple foreground/background PSO pairs. The effect was especially pronounced for pairs involving the face as a single foreground PSO ($p=0.0062$), face/face(a.) ($p<0.001$), face/tattoo ($p=0.0261$), and face/landmark ($p=0.0032$); where office settings consistently amplified PPL. In contrast, paper-based PSOs (e.g., birthdate, face(p.)) yielded lower overall comfortability levels but smaller differences between environments.
The violin plots in Figure 4 strengthen the analysis of H2 by visually illustrating the underlying distributional patterns. Across all eight statistically signifiant pairings, participants consistently reported lower comfortability level in the office setting compared to café, reaffirming the context sensitivity of perceived privacy. Figure 4 further reveals the changes in distribution spread, indicating a greater consensus about discomfort in office environments.
@@
Figure 4. Violin Plots of significant Mann-Whitney U test results for café and office environments (H2). Each subplot shows comfortability levels (1–5) for a foreground/background PSO pair; width reflects data density, and orange lines mark group mean differences.
These results suggest that not all PPLs of PSOs are equally shaped by the environmental context. Highly identifiable cues, such as faces, are perceived particularly private in professional settings, where co-presence may further heighten concerns. In contrast, more document-like PSOs show consistently low comfortability regardless of the environment. The violin plots add nuance to this interpretation by revealing how comfortability levels are lower in average in office settings, and more tightly distributed, indicating stronger agreement among participants. In conclusion, the context amplifies perceived privacy levels, and sensitivity is affected not only from single objects but also from their combinations and settings. This strengthens the claim that privacy-aware systems must move beyond binary recognition to incorporate compositional object definitions and contextual awareness.
4.4. Effect of Co-Presence (H3)
We evaluate H3 by testing whether the co-presence of background PSOs further increases the perceived privacy level (PPL) of a foreground PSO compared to the presence of a single background PSO. Since asking each participant to rate every possible combination would have made the survey longer and could lead to participant fatigue, we adopted a branching procedure: each participant first identified the background PSO that had the lowest effect on a given foreground PSO, and this “least effective” background PSO was paired with each remaining background PSO for the same foreground PSO. Table 6 reports how often each background PSO was assessed as the least effective, which determined the combinations evaluated.
Table 6. Background PSOs rated as having the least effect on comfortability level for each foreground PSO. Values indicate counts with percentages for Café ($n=51$) and Office ($n=41$). For each environment and foreground PSO, the background PSO most frequently selected is highlighted in bold.
Face Café / Office Misc. Paper Café / Office Computer/phone Screen Café / Office
Face (a.) 13 (26%) / 8 (20%) Face (p.) 15 (29%) / 6 (15%) Face (r.) 13 (26%) / 4 (10%)
Poster 9 (18%) / 10 (24%) Full Name 12 (24%) / 19 (46%) Date 36 (71%) / 32 (78%)
Medicine – / 2 (5%) Address 4 (8%) / 2 (5%) E-mail – / 1 (2%)
Tattoo 11 (22%) / 10 (24%) Birth Date 16 (31%) / 6 (15%) Username 2 (4%) / 2 (5%)
Landmark 18 (35%) / 11 (27%) Signature 4 (8%) / 8 (20%) Phone Number – / 2 (5%)
The resulting selection imbalance rendered formal statistical testing approaches inappropriate. Therefore, we report two complementary quantitative measures: (1) the mean change in the comfortability level when a second background PSO is added to the image and (2) the proportion of participants whose comfortability level decreased after this addition.
@@ 4.4. Effect of Co-Presence (H3)
Figure 5. Average comfortability level ratings for each foreground PSO, showing the baseline, the effect of the two background PSO most often rated as having the lowest effect when alone (see Table 6), and the effect of co-presence with both background PSOs present in the image (both).
Figure 5 illustrates the magnitude of mean changes when two least effective background PSOs (extracted from Table 6) are present in the image, labeled “both”. Table 7 reports the proportion of participants who decreased their level of comfortability when they observe two background PSOs. Both results converged to the same pattern: a small subset of PSO pairs consistently decreased the comfortability level (amplified perceived privacy level) beyond their single-object effects. For example, the co-presence of landmark/tattoo produced a considerable average drop in comfortability level ($0.66$) and in a substantial fraction of participants, the co-presence reduced the willingness to share. Pairs such as poster/tattoo and landmark/tattoo repeatedly ranked among the most impactful, with other notable combinations including e-mail/username and phone number/username on computer/phone screen. Most other co-presences showed minimal effects.
Table 7. Co-presence effects of background/background PSO pairs on perceived privacy levels for the foreground PSO (Face, Misc. paper, Screen). $n$ denotes the number of participants who rated the comfortability levels for a given pair. The % Decrease column shows the proportion of these participants who rated a lower comfortability level than images where each background PSO is shown separately, highlighting cases where co-presence amplifies privacy concerns.
@@
To further explore co-presence, we examined whether certain background PSOs exert dominance effect, meaning that when they appeared together with another background PSO, participant’s comfortability levels aligned almost exclusively with the more impactful cue. Table 8 summarizes these results. For the Face foreground PSO, medicine dominated almost two-thirds of the cases (65.3%), indicating that participants largely ignored the other background PSO where medicine was present. Similarly, for Misc. Paper, address (32.7%) and signature (32.0%) often disregarded the influence of other background PSOs. For the Screen, identifiers such as username (58.7%), phone number (53.1%), and e-mail (47.4%) showed strong dominance, while date (5.7%) rarely dominated other background PSOs. Importantly, lower $N$ values indicate that more participants initially considered these PSOs to be relatively more effective on their own, making their strong dominance effects even more notable. For example, medicine was often chosen as the most effective background PSO when in isolation and overwhelmingly dominated the participant’s judgement when paired with others.
Table 8. Dominance effects of background PSOs on foreground PSOs (Face, Misc. Paper, Screen) during co-presence with other background PSOs. Columns: Background PSO = background privacy sensitive object co-present with other background PSOs; N = number of co-presence ratings; D = number of dominance cases; %D = proportion of dominance cases (D/N).
@@
Demographic Effects:
Although not a primary focus of our study, we also examined whether demographic factors influenced privacy perceptions. Significant effects are summarized in Table 11 in Appendix A.3. Overall, gender effects were the most present, with female participants reporting higher comfortability levels, while other demographic factors showed more isolated differences. Our results suggest that privacy-aware systems may benefit from considering demographic variability, particularly when tailoring defaults for awareness mechanisms.
4.5. Qualitative Insights
The qualitative thematic analysis (refer to Section 3.3.2) of the responses to the open-ended questions from the survey yielded ten codes spanning across three discrete themes. The refined codebook is shown in Table 9. Our results indicate that participants have a latent cognitive model when self-evaluating the privacy of visual content that comprises multiple sensitive elements. In particular, the identified themes depict a layered model with heuristics, deeper reasoning, and consequence weighing as follows: First, the participants have mental shortcuts and norms (such as, “Everyone does it this way; So, it is fine!”) as comfort heuristics. If such heuristics raise a red flag, they dig deeper into inference (such as, “What if someone figures out more?”) and potential consequences (such as, “Would that be dangerous?”).
Table 9. Thematic analysis results.
@@ 4.5. Qualitative Insights
Inference and Linkability Contamination risks Concerns about being associated with unrelated or unwanted ideological affiliations that could create false reputation.
@@
Perceived Privacy Norms: This theme explores the comfort heuristics inferred by participants based on their interpretation of social expectations, consent norms, and perceived acceptability while sharing visual content online. Some example quotes from this theme are as follows.
@@
……….. if there’s someone else in the photo I might feel more comfortable sharing rather than being just me. ………..
……….. I usually post other humans when it is some social event, like a party or celebration, so by default everyone is aware pictures are taken and they usually say if they don’t like them so they don’t want them posted ………..
@@
The codes under this theme included instances where the participants mentioned the co-presence of other individuals and situations requiring mutual consent for the photo to be shared online. Furthermore, the codes also covered scenarios where the public availability of sensitive information reduced privacy concerns for revealing additional data. We excluded quotes related to non-human objects.
@@
Inference and Linkability: This theme captures deeper reasoning by a user about how information in an image can reveal hidden attributes or be linked across datasets to identify, profile, or deanonymize themselves or others present in the shared image. Some representative survey responses from this theme are quoted below.
@@
……….. Poster can contain certain propaganda that I do not affiliate myself with, so I wouldn’t want it showing up. ………..
……….. A combination of my (full) name plus all the other would be considered an information breach for me and I would not feel comfortable of people that I am not close with (as is the case in social media) knowing all this stuff about me plus it would create a high risk ………..
@@
The underlying codes of this theme capture scenarios of risks related to being associated with unrelated topics or the combination of multiple PII in the photo revealing more than intended. Furthermore, they also included concerns due to the disclosure of real-time locations directly through the cues in the shared photo or indirectly through the metadata. Some codes also reflect ad-hoc strategies employed by participants to mitigate the aforementioned risks.
@@
Irreversible Harms: This theme captures specific visual objects whose disclosure can cause long-lasting harms that cannot be easily remediated. The codes reflect instances where the participant expresses concerns about privacy risks associated with disclosure of, for example, distinctive physical attributes (e.g., tattoos or birthmarks) or health conditions. The codes also covered responses that mention well-known cyber-threat actors and the permanent consequences they could cause. General privacy concerns and non-sensitive information leaks were excluded from this theme.
@@
……….. Address is also very personal, especially to females, due to possibilities of stalkers, kidnappers, or robbers. Birthdate can be used for some type of fraud as well, given that in many countries the personal code contains birthday date. ………..
@@
In summary, the results from qualitative analysis emphasize various aspects of user mental models for visual privacy self-evaluation. The Perceived privacy norms theme reflects ad-hoc strategies for managing comfort. The Inference and linkability theme captures the risk assessment. The Irreversibility harms theme reflects on the potentially dangerous consequences if the risk is actualized.
5. Discussion
@@
One of the key observations from our work is that users’ privacy perceptions do not always translate into privacy behaviour or user actions. For example, the results from the privacy awareness questions (Table 3) show that the participants expressed general concern about online privacy, but not all examined the photos for personal information before sharing. Furthermore, the qualitative data from the survey displays several instances of privacy awareness among the participants, which appears to have been formed through public news media debates around privacy concerns and mandatory training at the workplace and educational institutions. Despite the awareness, users often fail to assess the privacy risks or control their behavior while sharing visual content on online platforms. The lack of proactive practices highlights that the current systems fail to bridge the gap between users’ mental models and actions.
@@
The complexity of privacy–utility tradeoffs further contributes to this gap. Usually, the foreground privacy-sensitive objects (PSOs) carry inherent sensitivity. Still, people may choose to share them due to the tangible benefits they provide, such as identity verification, professional collaboration, or social connection. Our results show that the perceived privacy risks associated with these PSOs are not fixed, but dynamically reshaped by the presence of background PSOs. For example, a selfie becomes more sensitive when medicine is visible, a document becomes more risky to share when a signature is present, and a screenshot becomes problematic when the username or phone number is exposed. Such contextual amplification indicates that privacy–utility tradeoffs, especially in the case of visual contents, are a rather complicated process beyond a simple binary choice. The complexity in privacy decision-making arises from users’ attempts to weigh multiple contextual factors and engage in multidimensional self-negotiation for the tradeoffs.
@@
These observations emphasize the need for novel technical interventions and research opportunities to explore the notion of context-aware privacy. In this realm, the hypothesis evaluation and results from our study provide insights into design principles and required characteristics of a context-aware privacy system. From a system design point of view, such a system should avoid uniform treatment of visual elements or objects and adapt privacy sensitivity based on contextual cues. Moreover, the design should consider different types of contextual cues in the visual content. One of the types includes the subtler and latent cues captured in the visual scenes, such as religious symbols, political expressions, or social affiliations. These are not easily identifiable as discrete objects but are often regarded as very sensitive. The other type of contextual cues are visual elements (e.g., medicine) that have a dominance effect, overriding other cues. The dominant cues disproportionately influence user perceptions, as users often anchor their privacy judgments on the most dominant cue rather than considering the full context. Moreover, we also observed that the co-presence of certain visual elements can lead to even higher perceived privacy risks. Towards this end, we argue that privacy-aware system design benefits from integrating object-level detection with contextual inference while also accounting for the dominance and co-presence of visual elements.
@@
From a HCI design perspective, privacy-aware systems should provide automated assistance and fine-grained control to users. The assistance can be provided in the form of nudges, via visual highlighting and prompts, that guide users towards privacy-oriented actions without requiring extra effort. Alternatively, assistance can also be provided through the automatic detection of a highly sensitive object or contextual cue in the background that could amplify the privacy perception. Furthermore, the automated assistance should be suggesting appropriate privacy controls. The controls should be designed in such a way that the users can selectively obfuscate specific sensitive objects of the image in a way that it still retains utility and maintains users’ privacy when shared in the redacted form. In this realm, we argue that a human-centered design approach can assist users in translating their privacy perceptions into concrete actions, reducing cognitive load and improving decision-making.
@@
The existing online social media and instant communication platforms, which deal with user-generated visual content, largely lack user assistance and nuanced controls for privacy. The security and privacy settings for the user accounts offered by some of these platforms fail to address the privacy risks highlighted in our work. The results and discussions from our work can provide actionable insights to integrate visual privacy protection components. These insights can also guide usable system design themes of research works and development of future technologies that deliver personalized privacy protection.
5.1. Limitations and Future Work

### v2

@@ IV-C Effect of Environment (H2)
To evaluate H2, we first compared the average comfortability levels between environments. The comfortability level was consistently lower in office settings than in café, with the strongest drop observed for Face as a foreground PSO (Face: 0.67-0.67; Misc. Paper: 0.37-0.37; Screen: 0.31-0.31). To systematically assess these differences, we applied the Mann–Whitney U test. As shown in Table IV, participants reported significantly lower comfortability (i.e., higher perceived privacy level, PPL) in office settings across multiple foreground/background PSO pairs. The effect was especially pronounced for pairs that involve the face as a single foreground PSO (p=0.0062p\!=\!0.0062), face/face(a.) (p<0.001p\!<\!0.001), face/tattoo (p=0.0261p\!=\!0.0261), and face/landmark (p=0.0032p\!=\!0.0032); where office settings consistently amplified PPL. In contrast, paper-based PSOs (e.g., birthdate, face(p.)) yielded lower overall comfortability levels but smaller differences between environments. The violin plots in Appendix Figure 8 strengthen the analysis of H2 by visually illustrating the underlying distributional patterns. Across all eight statistically significant pairings, participants consistently reported lower comfortability level in the office setting compared to café, reaffirming the context sensitivity of perceived privacy. Figure 8 further reveals changes in distribution spread, indicating a greater consensus on discomfort in office environments.
@@
Figure 4: Average comfortability level ratings for each foreground PSO, showing the baseline, the effect of the two background PSO most often rated as having the lowest effect when alone (see Table V), and the effect of co-presence with both background PSOs present in the image (both).
TABLE V: Background PSOs rated as having the least effect on comfortability level for each foreground PSO. Values indicate counts with percentages for Café ($n=51$) and Office ($n=41$). For each environment and foreground PSO, the background PSO most frequently selected is highlighted in bold and blue.
Face Café / Office Misc. Paper Café / Office Screen Café / Office Face (a.) 13 (26%) / 8 (20%) Face (p.) 15 (29%) / 6 (15%) Face (r.) 13 (26%) / 4 (10%) Poster 9 (18%) / 10 (24%) Full Name 12 (24%) / 19 (46%) Date 36 (71%) / 32 (78%) Medicine – / 2 (5%) Address 4 (8%) / 2 (5%) E-mail – / 1 (2%) Tattoo 11 (22%) / 10 (24%) Birth Date 16 (31%) / 6 (15%) Username 2 (4%) / 2 (5%) Landmark 18 (35%) / 11 (27%) Signature 4 (8%) / 8 (20%) Phone Number – / 2 (5%)
These results suggest that not all PPLs of PSOs are equally shaped by the environmental context. Highly identifiable cues, such as faces, are perceived particularly private in professional settings, where co-presence may further heighten concerns. In contrast, more document-like PSOs show consistently low comfortability regardless of the environment. The violin plots add nuance to this interpretation by revealing how comfortability levels are lower in average in office settings, and more tightly distributed, indicating stronger agreement among participants. In conclusion, the context amplifies perceived privacy levels, and sensitivity is affected not only by single objects but also by their combinations and settings. This strengthens the claim that privacy-aware systems must move beyond binary recognition to incorporate compositional object definitions and contextual awareness.
IV-D Effect of Co-Presence (H3)
TABLE VI: Co-presence effects of background/background PSO pairs on perceived privacy levels for the foreground PSO (Face, Misc. Paper, Screen). $n$ denotes the number of participants who rated the comfortability levels for a given pair. The % Decrease column shows the proportion of these participants who rated a lower comfortability level than images where each background PSO is shown separately, highlighting cases where co-presence amplifies privacy concerns.
@@ IV-D Effect of Co-Presence (H3)
TABLE VII: Dominance effects of background PSOs on foreground PSOs (Face, Misc. Paper, Screen) during co-presence with other background PSOs. Columns: Background PSO = background privacy sensitive object co-present with other background PSOs; N = number of co-presence ratings; D = number of dominance cases; %D = proportion of dominance cases (D/N).
@@
We evaluated H3 by testing whether the co-presence of background PSOs further increases the perceived privacy level (PPL) of a foreground PSO compared to the presence of a single background PSO. To avoid survey fatigue from rating all possible combinations, we adopted a branching procedure: each participant first identified the background PSO that had the lowest effect on a given foreground PSO, and this “least effective” background PSO was paired with each remaining background PSO for the same foreground PSO. Table V reports how often each background PSO was assessed as the least effective, determining the combinations evaluated. The resulting selection imbalance made formal statistical testing approaches inappropriate. Therefore, we report two complementary quantitative measures: (1) the mean change in the comfortability level when a second background PSO is added to the image and (2) the proportion of participants whose comfortability level decreased after this addition.
Figure 4 illustrates the magnitude of mean changes when two least effective background PSOs (extracted from Table V) are present in the image, labeled “both”. Table VI reports the proportion of participants who decreased their level of comfortability when they observe two background PSOs. Both results converged to the same pattern: a small subset of PSO pairs consistently decreased the comfortability level (amplified perceived privacy level) beyond their single-object effects. For example, the co-presence of landmark/tattoo produced a considerable average drop in comfortability level ($0.66$) and in a substantial fraction of the participants, the co-presence reduced the willingness to share the foreground PSO. Pairs such as poster/tattoo and landmark/tattoo repeatedly ranked among the most impactful, with other notable combinations including e-mail/username and phone number/username on computer/phone screen. Most other co-presences showed minimal effects.
To further explore co-presence, we examined whether certain background PSO exerts dominance effect, meaning that when they appeared together with another background PSO, participant’s comfortability levels aligned almost exclusively with the more influential cue. Table VII summarizes these results. For the Face foreground PSO, medicine dominated almost two-thirds of the cases (65.3%), indicating that the participants largely ignored the other background PSO where medicine was present. Similarly, for Misc. Paper, address (32.7%) and signature (32.0%) often dominated the influence of other background PSOs. For the Screen, identifiers such as username (58.7%), phone number (53.1%), and e-mail (47.4%) showed strong dominance, while date (5.7%) rarely dominated other background PSOs. Importantly, lower $N$ values indicate that more participants initially considered these PSOs to be relatively more effective on their own, making their strong dominance effects even more notable. For example, medicine was often chosen as the most effective background PSO when in isolation and overwhelmingly dominated the participant’s judgment when paired with others.
@@
Demographic Effects: We also examined whether demographic factors influenced privacy perceptions. Significant effects are summarized in Table XI in Appendix A-C. In general, gender effects were the most present, with female participants reporting higher comfortability levels, while other demographic factors showed more isolated differences. Our results suggest that privacy-aware systems may benefit from considering demographic variability, particularly when tailoring defaults for awareness mechanisms.
IV-E Qualitative Insights
Qualitative thematic analysis (refer to Section III-D2) of responses to open-ended survey questions yielded ten codes spanning across three discrete themes. The refined codebook is shown in Table VIII. Our results indicate that participants have a latent cognitive model when self-evaluating the privacy of visual content that comprises multiple sensitive elements. In particular, the identified themes depict a layered model with heuristics, deeper reasoning, and consequence weighing as follows: First, the participants have mental shortcuts and norms (such as “Everyone does it this way; So, it is fine!”) as comfort heuristics. If such heuristics raise a red flag, they dig deeper into inference (such as “What if someone figures out more?”) and potential consequences (such as “Would that be dangerous?”).
TABLE VIII: Thematic analysis results.
@@ IV-E Qualitative Insights
Inference & Linkability Contamination risks Concerns about being associated with unrelated or unwanted ideological affiliations that could create false reputation.
@@
Perceived Privacy Norms: (See quotes Q1 and Q2) This theme explores the comfort heuristics inferred by participants based on their interpretation of social expectations, consent norms, and perceived acceptability while sharing photos online.
@@
Q1: ……….. if there’s someone else in the photo I might feel more comfortable sharing rather than being just me. ………..
@@
This theme’s codes included instances where the participants mentioned the co-presence of other individuals and situations requiring mutual consent for the photo to be shared online. The codes also covered scenarios where the public availability of sensitive information reduced privacy concerns for revealing additional data. We excluded quotes related to non-human objects.
Q2: ………. I usually post other humans when it is some social event, like a party or celebration, so by default everyone is aware pictures are taken and they usually say if they don’t like them so they don’t want them posted ………..
@@
Inference and Linkability: (See quotes Q3 and Q4) This theme captures a deeper reasoning by a user about how information in an image can reveal hidden attributes or be linked across datasets to identify, profile, or de-anonymize themselves or others present in the shared image.
@@
Q3:….. Poster can contain certain propaganda that I do not affiliate myself with, so I wouldn’t want it showing up. ……
@@
The underlying codes of this theme capture scenarios of risks related to being associated with unrelated topics or the combination of multiple PII in the photo that reveal more than intended. Furthermore, they also included concerns about the disclosure of real-time locations directly through the cues in the shared photo or indirectly through the metadata. Some codes also reflect ad-hoc strategies employed by participants to mitigate the aforementioned risks.
Q4:………. A combination of my (full) name plus all the other would be considered an information breach for me and I would not feel comfortable of people that I am not close with (as is the case in social media) knowing all this stuff about me plus it would create a high risk ………..
@@
Irreversible Harms: (See quote Q5) This theme captures specific visual objects whose disclosure can cause long-lasting harms that cannot be easily remediated. The codes reflect instances where the participant expresses concerns about privacy risks associated with disclosure of, for example, distinctive physical attributes (e.g., tattoos or birthmarks) or health conditions. The codes also covered responses that mention well-known cyber-threat actors and the permanent consequences they could cause. General privacy concerns and non-sensitive information leaks were excluded from this theme.
@@
Q5:….. Address is also very personal, especially to females, due to possibilities of stalkers, kidnappers, or robbers. Birthdate can be used for some type of fraud as well, given that in many countries the personal code contains birthday date……
@@
In summary, the results of the qualitative analysis emphasize various aspects of user mental models for visual privacy self-evaluation. The Perceived Privacy Norms theme reflects ad-hoc strategies for managing comfort. Inference and Linkability captures the risk assessment. The Irreversibility Harms theme reflects on the potential dangers if the risk is actualized.
V Discussion
@@
One of our key observations is that users’ privacy perceptions do not always translate into privacy behavior or user actions. For example, the results from the privacy awareness questions (Table II) show that the participants expressed general concern about online privacy, but not all examined the photos for personal information before sharing. Qualitative data from the survey also displays several instances of privacy awareness among participants, which appears to have been formed through public news media debates around privacy concerns and mandatory training at the workplace and educational institutions. Despite the awareness, users often fail to assess the privacy risks or control their behavior while sharing visual content on online platforms [cheung2016evaluating, barnes2006privacy, gerber2018explaining]. The lack of proactive practices highlights that current systems fail to bridge the gap between users’ mental models and actions.
@@
The complexity of privacy–utility tradeoffs further contributes to this gap. Usually, foreground PSOs carry inherent sensitivity. However, people may choose to share them because of the tangible benefits they provide, such as identity verification, professional collaboration, or social connection. Our results show that the perceived privacy risks associated with these PSOs are not fixed but dynamically reshaped by the presence of background PSOs. For example, a selfie becomes more sensitive when medicine is visible, a document becomes more risky to share when a signature is present, and a screenshot becomes problematic when the username or phone number is exposed. Such contextual amplification indicates that privacy–utility tradeoffs, especially in the case of visual contents, are a rather complicated process beyond a simple binary choice. The complexity in privacy decision-making arises from users’ attempts to weigh multiple contextual factors and engage in multidimensional self-negotiation for the tradeoffs.
@@
These observations emphasize the need for novel technical interventions and research opportunities to explore the notion of context-aware privacy. In this realm, the hypothesis evaluation and results from our study provide insight into the design principles and required characteristics of a context-aware privacy system. From a system design point of view, such a system should avoid uniform treatment of visual elements and adapt privacy sensitivity based on contextual cues. Also, the design should consider two broad categories of contextual cues: (i) subtle, latent cues (e.g., religious symbols or political affiliations) that are hard to detect, and (ii) categorically sensitive elements (e.g., medicine) that have a dominance effect by overriding other cues disproportionately influencing user perceptions, as users often base their privacy judgments on such cues instead of the full context. Moreover, we observed that the co-presence of certain visual elements can lead to even higher perceived privacy risks. Thus, we argue that privacy-aware system design benefits from integrating object-level detection with contextual inference that accounts for the dominance and co-presence of visual elements.
@@
From an HCI design perspective, privacy-aware systems should provide automated assistance and fine-grained control to users. The assistance can be provided in the form of nudges, via visual highlighting and prompts, that guide users towards low-burden, privacy-oriented actions. Alternatively, assistance can also be provided through the automatic detection of highly sensitive objects or contextual cues in the background that could amplify the perception of privacy. Automatic assistance should also suggest appropriate privacy controls, allowing users to selectively obfuscate specific sensitive objects of the image while retaining the utility of the image and preserving privacy in its redacted form. Within this context, we argue that a human-centered design approach can assist users in translating their privacy perceptions into concrete actions, reducing cognitive load and improving decision-making.
@@
Existing online social media and instant communication platforms, which deal with user-generated visual content, largely lack user assistance and nuanced controls for privacy. The privacy settings for the user accounts offered by some of these platforms fail to address the risks highlighted in our work. The results and discussions of our work can provide actionable insights to integrate visual privacy protection components. These insights can also guide usable system design themes of research works and development of future technologies that deliver personalized privacy protection.
V-A Limitations and Future Work